# HeadGaS: Real-Time Animatable Head Avatars via 3D Gaussian Splatting

### v1

@@ 3.2 Feature Blending Formulation
Here we describe how we extend the 3DGS representation with animation capabilities. The vanilla 3DGS model does not inherently allow for this, as it learns a static set of parameters, which is the same for all frames. Inspired by 3DMMs, the goal of our model is to explore a blending mechanism for the 3D Gaussian components, using the pre-computed expression parameters as blending weights. Namely, we want each Gaussian to change color and opacity based on the current expression $i$. This leads to 3D Gaussians with dynamic appearance which occasionally appear and vanish depending on the current expression, and additionally allow color changes for non-rigid appearance effects such as wrinkles. For instance, for a certain expression vector in which the subject exhibits neutral expression with closed mouth, we want the Gaussians corresponding to lip regions to be opaque, while if the person starts to laugh, the lips become transparent and reveal the teeth regions.
The naive approach of extending a 3DGS representation to such dynamic appearance, would be to have a base of each Gaussian parameter, one for each Gaussian, and blend them based on the expression weights. However, as these values have an explicit meaning (i.e. color, position), a multiplication with expression weights that are not even learnable, makes this formulation limiting and prone to artifacts, as we also show in our experiments. Therefore, we extend every 3D Gaussian with a base of latent features $\bm{F}\in\mathbb{R}^{B\times f_{dim}}$, Figure 2. Our animatable 3D Gaussian representation then becomes $\mathcal{G}_{a}=(\bm{\Sigma},\bm{\mu},\bm{F})$.
The latent base is optimized together with the other parameters of the 3D Gaussian. At each iteration, we leverage the respective expression weights $\bm{e}_{i}\in\mathbb{R}^{B}$ of the current frame $i$, to blend the feature base $\bm{F}$ into a 1D vector $\bm{f}_{i}\in\mathbb{R}^{f_{dim}}$
@@
where $\psi$ denotes cosine positional encoding.
Thereby the learned color is a 1D vector $\bm{c}_{i}\in\mathbb{R}^{3(k+1)^{2}}$, while the learned opacity is a scalar $\alpha_{i}\in\mathbb{R}$. Our MLP is composed of two linear layers followed by LeakyReLU activation [41], where the hidden layer has $64$ channels. The last layer consists of two branches, for namely color and opacity prediction. We use a sigmoid activation function at the end of the opacity branch to constrain it to be in its natural range $[0,1]$. As color and opacity are learned via the MLP, we omit them from the explicit Gaussian parameters.
Interestingly, though changing the centers and rotations instead, would be quite an intuitive alternative when it comes to 3D Gaussian animation, the proposed approach results in much better performance (see Sec. 4.2).
@@ 3.3 Rendering
To render frame $i$, we employ the respective expression $\bm{e}_{i}$ to populate each Gaussian with expression-dependent color and opacity. Then, the Gaussians are rendered using the camera view $\bm{W}_{i}$. Thereby, following [17] we employ the splatting technique [44]. Given a viewing transform $\bm{W}$ as well as the Jacobian of the affine approximation of the projective transformation $\bm{J}$, the covariance matrix $\bm{\Sigma}^{\prime}$ in camera coordinates can be obtained from
@@
The Gaussian splats are finally rendered via a tile-based differentiable rasterizer [17] that pre-sorts all primitives of an image at once.
@@ 3.4 Optimization
We initialize the 3D Gaussians centers with 2500 points. Whenever available, these points are a subset of vertices from the tracked 3DMM meshes (e.g. FLAME based data released by [48, 47]). As there is no mesh available for the data from [11], we sample random points within the provided near and far bounds. Empirically we found that initializing the latent features $\bm{F}$ with zeros led to the most stable solution. The model is optimized by rendering the learned Gaussians and comparing the resulting image $I_{\text{r}}$ against the ground truth $I_{\text{gt}}$. We minimize the following loss objective
$\mathcal{L}_{\text{total}}=\mathcal{L}_{1}(I_{\text{r}},I_{\text{gt}})+\lambda_{s}\mathcal{L}_{\text{SSIM}}(I_{\text{r}},I_{\text{gt}})+\lambda_{p}\mathcal{L}_{p}(I_{\text{r}},I_{\text{gt}})$ (6)
@@
HeadGaS (Ours) 0.0039 24.79 0.889 0.117 0.004
Adaptive densification and pruning
3.5 Implementation details
Our learning rates for the MLP $\phi(\cdot)$, positions $\mu$, latent features $\bm{F}$, scale $S$ and rotation $R$ are namely $0.00016$, $0.00016$, $0.0025$, $0.005$ and $0.001$. We set the latent feature dimensionality to $f_{dim}=32$. All Gaussians are fed as a single batch into the MLP. The $\mathcal{L}_{p}$ loss is based on a VGG network [34] and has a weight of $\lambda_{p}=0.1$, while $\lambda_{s}=0.2$. We activate the $\mathcal{L}_{p}$ loss after $10k$ iterations such that it does not conflict with photometric loss at the early stage of learning. The densification starts after 500 iterations and stops with 15k iterations. In our experiments we use an SH degree of $k=3$. We train our models on a single Tesla $V100$ GPU for $50k$ iterations taking about 1 hour.
4 Experiments
In this section we describe the evaluation protocol and the experiments carried out to evaluate the effectiveness of the proposed method. Then, we present results in three different scenarios, such as same-person pose and expression driving, novel view synthesis as well as cross-subject expression driving.
Figure 3: Qualitative evaluation comparing the proposed model against INSTA [48], PointAvatar [47] and NeRFBlendShape [11] baselines, namely on the a) INSTA data, b) NBS data and c) PointAvatar data. We provide close-ups on the right of each example, to highlight our method’s ability to capture details like teeth, wrinkles and reflections.
We evaluate our model on multiple datasets, made publicly available by recent works such as NeRFBlendShape (we dub this NBS data), INSTA, and PointAvatar. The NBS data [11] contains a set of monocular videos from 8 subjects, where the last $500$ frames in each subject constitute the test set. The INSTA dataset [48] contains 10 subjects, where the last $350$ frames of each sequence are kept as test set. We additionally evaluate on the 3 subjects made available by PointAvatar [47], where the test sets contain between $880-1800$ frames per subjects. For fairness of comparison, when training our method in all three datasets, we utilize the tracking data (head poses and expression weights) provided originally by the authors, such that tracking quality does not affect the comparison. We use exactly the same split as the original methods. We use a subset of about $1500-2500$ frames from the training set of each subject to train our model.
To assess the quality of the synthesized images we report common metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) [45] and the Mean Squared Error (L2). Additionally, we report time measurements for the rendering in seconds. We compare against common baselines such as NeRFBlendShape [11], INSTA [48], PointAvatar [47], NHA [13], IMAvatar [46] and NeRFACE [10]. The evaluation is carried out using the respective official code repositories, as well as their official checkpoints whenever available.
@@ 4.1 Same-subject novel expression driving
In this experiment we drive the learned subjects using the novel poses and expression parameters of their test set. Table 1 reports the results of the metric comparison against baselines. The three different blocks show namely results on the data released by INSTA, NBS and PointAvatar. Figure 3 illustrates the qualitative comparison with the most recent baselines in all three datasets, namely a) INSTA data, b) NBS data and c) PointAvatar data. We observe that the proposed method outperforms all baselines on the INSTA and NBS datasets in all metrics, by namely about $1dB$ and $2dB$. Referring to Figure 3, we observe that HeadGaS leads to higher fidelity to ground truth, less artifacts, and identity preservation for all subjects. Interestingly, as INSTA relies on mesh deformations, it exhibits artifacts such as noticeable triangles (Figure 3 a), INSTA). Moreover, our model preserves better details, such as wrinkles, eyebrows, teeth and glass reflections, while other baselines [47, 48, 11] often struggle in such aspects. The comparison against the PointAvatar baseline shows that we are superior on the INSTA data by about $1dB$. On the PointAvatar dataset, our method surpasses the baseline in terms of LPIPS and SSIM, while having lower PSNR ($1.2$ dB). Overall, looking at both datasets, HeadGaS is superior to PointAvatar. Also qualitatively, we can see that PointAvatar results have distortions of some parts that undergo significant transformation (Figure 3 c)) or inaccurate teeth regions (Figure 3 b)). We believe that, as an explicit method, PointAvatar generalizes better for under-observed expressions (resulting in slightly higher PSNR). However, as these cases struggle with deformation realism, the structural metrics are worse and the problem remains unsolved. Finally, as the table shows, we surpass the rendering time of all baselines for 512 resolution by at least a factor of 10. We refer the reader to the supplementary material for more qualitative results and videos.
@@ 4.2 Ablations
In this section we ablate the HeadGaS components. First, we train a model that does not use learned features for the blending, but rather a base of colors and opacities, and uses the expression weights to obtain the final color and opacity as a weighted average. We refer to this model as Ours w/o MLP. In addition, since an intuitive alternative for dynamic Gaussians would be to deform the points (rather than adapting color and opacity) we introduce Ours w/ $\Delta(\mu,R)$, which uses the learned feature basis (and a similar MLP) to rather shift the positions $\mu$ and transform the rotations $R$. Further, we run a model without the perceptual loss, i.e. Ours w/o $\mathcal{L}_{p}$, to ablate its effectiveness. Additionally, to confirm the contribution of using the expression parameters as a weight for blending Gaussian features, rather than a simple condition to the MLP, we ablate a variant named Ours w/o blending. More details can be found in the supplement. Table 2 reports the quantitative evaluation of our model components. We observe that using the expression parameters as a simple condition (Ours w/o blending) leads to noticeably inferior performance. Also, applying a shift to the positions and transformation to the rotations leads to worse results (Ours w/ $\Delta(\mu,R)$). We believe this is because transforming points is a more complex problem to optimize. Figure 4 confirms these results and reveals floater artifacts and inaccurate expressions. Further, blending the explicit parameters directly (Ours w/o MLP) leads to a a worse performance than our neural variant. Though the PSNR gap is not very major, we notice a drastic visual effect on the highly dynamic areas, relevant to the blending, as illustrated in Figure 4. Finally, we see that adding a perceptual loss term $\mathcal{L}_{p}$ leads to an improvement in most metrics.
@@
Table 2: Ablation of HeadGaS components on the INSTA dataset.
Figure 4: Qualitative ablation on the INSTA data illustrating the effect of MLP and the Gaussian deformation alternative.
@@ 4.3 Novel view synthesis
In Figure 5 we show our generated avatars from multiple views, including the original test set camera (left) and two additional viewpoints (right). Thereby we render the same facial expression. We observe that our model can deliver expressions that are consistent across different views.
Figure 5: Qualitative HeadGaS results on multiple camera views for the same expression parameters.
@@ 4.4 Cross-subject expression driving
So far we have shown that our model is capable of generalizing to novel expression parameters corresponding to the same person, in time-frames not observed during training. Figure 6 complements these results with cross-subject driving, i.e. use the facial expression from another (ground truth) source subject to drive a target subject. As can be seen in the figure, our model is capable of transferring expression across different subjects at a reasonable quality.
Figure 6: Cross-subject expression driving.
Figure 7: Limitations. Left: subject only observed in neutral expression in non-frontal view. Right: Camera view that is far from the training observations.
@@ 5 Limitations and ethical consideration
As a method that relies on tracked data, i.e. head poses and expressions, HeadGaS is affected by failures of the head tracker. This can take the form of inaccurate facial expressions, or blurriness in case of pose inaccuracy in the training data (Figure 7). Further, as a data driven method, HeadGaS requires a reasonable coverage of expressions spread among different views. For instance, if a head changes pose only in neutral expression, and diverse expressions are captured only in frontal view, it would be difficult to capture a non-neutral expression from side view (Figure 7, left).
We carried out this work with subjects which are publicly available [48, 11, 47]. They are a combination of lab captures and public YouTube videos. We have not captured additional subjects. The authors claim that the subjects have given their consent for using their data for research purposes. The goal of this work is to immerse user experience in AR/VR applications, in which they have full control of their data and animation. However, we acknowledge that our method could be extended to generate fake content of people, without their consent, with the purpose to spread false information. We strictly oppose such behaviour.

### v2

@@ 3.2 Feature Blending Formulation
Here we describe how we extend the 3DGS representation with animation capabilities. The vanilla 3DGS model does not inherently allow for this, as it learns a static set of parameters, which is the same for all frames. Inspired by 3DMMs, the goal of our model is to explore a blending mechanism for the 3D Gaussian components, using the pre-computed expression parameters as blending weights. Namely, we want each Gaussian to change color and opacity based on the current expression $i$. This leads to 3D Gaussians with dynamic appearance which occasionally appear and vanish depending on the current expression, and additionally allow color changes for non-rigid appearance effects such as wrinkles. For instance, referring to Figure 2, Gaussians of closed lips visible in frame $i$ will turn transparent in frame $j$, as the jaw opens, while another set of Gaussians at a different location will become visible to render open lips. The model will thus learn multiple Gaussians corresponding to the same region in the face, such that these can become opaque as needed.
With the goal of enabling such dynamic appearance, we extend every 3D Gaussian with a basis of latent features $\bm{F}\in\mathbb{R}^{B\times f_{dim}}$, (see Figure 3). Our animatable 3D Gaussian representation then becomes $\mathcal{G}_{a}=(\bm{\Sigma},\bm{\mu},\bm{F})$. The latent base is optimized together with the other parameters of the 3D Gaussian. At each iteration, we leverage the respective expression weights $\bm{e}_{i}\in\mathbb{R}^{B}$ of the current frame $i$, to blend the feature basis $\bm{F}$ into a 1D vector $\bm{f}_{i}\in\mathbb{R}^{f_{dim}}$
@@
where $\psi$ denotes sinusoidal positional encoding, the learned color is a 1D vector $\bm{c}_{i}\in\mathbb{R}^{3(k+1)^{2}}$, and the learned opacity is a scalar $\alpha_{i}\in\mathbb{R}$. As most of the dynamic effects are already captured by the per-Gaussian feature bases $\bm{F}$, we are able to use a very small MLP that does not compromise the rendering speed. Our MLP is composed of only two linear layers, each followed by LeakyReLU activation [53], where the hidden layer has $64$ channels. The last layer consists of two branches, i.e. for color and opacity prediction. We use a sigmoid activation function at the end of the opacity branch to constrain it to be in its appropriate range $[0,1]$. As color and opacity are learned via the MLP, we omit them from the explicit optimizable Gaussian parameters.
An alternative to blending in the latent space, would be to directly define a basis of explicit Gaussian parameters and similarly blend them based on the expression weights. However, as these values have an explicit meaning (i.e. color, position), a multiplication with expression weights that are not even learnable, makes this formulation limiting and prone to artifacts, as we also show in our ablation (Sec. 4.2, Ours w/o MLP). Interestingly, even though changing the centers and rotations of the Gaussian splats (instead of modifying opacity and colour) would be an intuitive mechanism when it comes to 3D Gaussian animation, our proposed approach results in much better performance (see Sec. 4.2, Ours w/ $\Delta(\mu,R)$). We additionally show that the proposed feature blending strategy is superior to the straight-forward approach of using the expression vector as a condition to the MLP (Ours w/o blending).
@@ 3.3 Rendering
To render frame $i$, we employ the respective expression $\bm{e}_{i}$ to populate each Gaussian with expression-dependent color and opacity. Then, the Gaussians are rendered using the camera view $\bm{W}_{i}$, and similarly to Kerbl et al. [20] we perform the splatting technique on the primitives [57]. Given a viewing transform $\bm{W}$ as well as the Jacobian of the affine approximation of the projective transformation $\bm{J}$, the covariance matrix $\bm{\Sigma}^{\prime}$ in camera coordinates can be obtained from
@@
The Gaussian splats are then rendered via a tile-based differentiable rasterizer [20] that pre-sorts all primitives of an image at once.
@@ 3.4 Optimization
We initialize the 3D Gaussians centers with 2500 points. Whenever available, these points are a subset of vertices from the tracked 3DMM meshes (e.g. FLAME based data released by prior works [62, 61]). As there is no mesh available for the data from Gao et al. [14], we sample random points within the given near and far bounds. Empirically we found that initializing the latent features $\bm{F}$ with zeros led to the most stable solution. The model is optimized by rendering the learned Gaussians and comparing the resulting image $I_{\text{r}}$ against the ground truth $I_{\text{gt}}$. We minimize the following loss objective
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{1}(I_{\text{r}},I_{\text{gt}})+\lambda_{s}\mathcal{L}_{\text{SSIM}}(I_{\text{r}},I_{\text{gt}})+\lambda_{p}\mathcal{L}_{p}(I_{\text{r}},I_{\text{gt}})$ (6)
@@
3.4.1 Adaptive densification and pruning
3.5 Implementation Details
The learning rates for the MLP $\phi(\cdot)$, positions $\mu$, latent features $\bm{F}$, scale $S$ and rotation $R$ are namely $1.6{\cdot}10^{\text{-}4}$, $1.6{\cdot}10^{\text{-}4}$, $0.0025$, $0.005$ and $0.001$. We set the latent feature dimensionality to $f_{dim}{=}32$. For the FLAME tracking [62] we only use the first 52 expression weights, i.e. $B{=}52$. All Gaussians are fed as a single batch into the MLP. The $\mathcal{L}_{p}$ loss is based on a VGG network [44] and has a weight of $\lambda_{p}{=}0.1$, while $\lambda_{1}{=}0.8$ and $\lambda_{s}{=}0.2$. We activate the $\mathcal{L}_{p}$ loss after $10k$ iterations such that it does not conflict with photometric loss at the early stage of learning. To save computation, we apply $\mathcal{L}_{p}$ on the image region defined by the head bounding box. The densification starts after 500 iterations and stops with 15k iterations. In our experiments we use an SH degree of $k{=}3$. We train our models on one Tesla $V100$ GPU for $50k$ iterations taking about 1 hour.
4 Experiments
In this section we describe the evaluation protocol followed by quantitative and qualitative results in three different scenarios, e.g. same-subject novel expression and novel view rendering, as well as cross-subject expression driving.
@@ 4 Experiments
HeadGaS (Ours) 0.0029 25.99 0.897 0.108 0.004
We evaluate our model on three datasets, made publicly available by recent works such as NeRFBlendShape (we dub this NBS data), INSTA, and PointAvatar. The NBS data [14] contains a set of monocular videos from 8 subjects, where the last $500$ frames in each subject constitute the test set. The INSTA dataset [62] contains 10 subjects, with the last $350$ frames of each sequence being the test set. We additionally evaluate on the 3 subjects made available by PointAvatar [61], where the test sets contain between $880-1800$ frames per subject. For fairness of comparison, when training our method in all three datasets, we use the same splits and utilize the tracked data (head poses and expression weights) provided originally by the authors, such that tracking quality does not affect the comparison. To train our model we use a subset of about $1500-2500$ frames from the training set of each subject. To assess the quality of the synthesized images we report common metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) [59] and the Mean Squared Error (L2). All metrics are computed using white background for non-face regions. Further, we report rendering times in seconds. We compare against common baselines such as NeRFBlendShape [14], INSTA [62], PointAvatar [61], NHA [16], IMAvatar [60], AvatarMAV [55], FLARE [4] and NeRFACE [13]. The evaluation is carried out using their official code repositories, as well as their official checkpoints whenever available.
@@ 4.1 Same-Subject Novel Expression Driving
Table 1 reports the results of the metric comparison against baselines on the respective test sets. The three different blocks show namely results on the data released by INSTA, NBS and PointAvatar. Figure 4 illustrates the qualitative comparison with the most recent baselines in all three datasets. We observe that the proposed method outperforms all baselines on the INSTA and NBS datasets in all metrics, with a PSNR gap of about $2$ dB. Referring to Figure 4, we observe that HeadGaS leads to higher fidelity to ground truth, less artifacts, and identity preservation for all subjects. Interestingly, as INSTA relies on mesh deformations, it exhibits artifacts such as noticeable triangles on the skin surface (Figure 4a). Moreover, our model preserves better details, such as facial expressions, wrinkles, eyebrows, teeth and glass reflections, while other baselines [61, 62, 14] often struggle in such aspects. The comparison against the PointAvatar baseline shows that we are superior on the INSTA data by about $2$ dB. On the 3 subjects of the PointAvatar dataset, our method surpasses the baseline in terms of LPIPS and SSIM, while having very similar PSNR. Overall, looking at both datasets, the performance of HeadGaS is superior to that of PointAvatar. Also qualitatively, we can see that PointAvatar results have distortions of some parts that undergo significant transformation, including inaccurate teeth deformation (Figure 4c). We believe that, as an explicit method, PointAvatar generalizes well for under-observed expressions (resulting in comparable PSNR). However, as these cases struggle with deformation realism (teeth wrongly deform in the same way as the mouth), the structural metrics are worse. Finally, as the table shows, we improve the rendering time of all baselines for 512 resolution by at least a factor of 10. We refer the reader to the supplement for more qualitative results and videos.
Figure 4: Qualitative evaluation comparing the proposed model against INSTA [62], PointAvatar [61] and NeRFBlendShape [14] baselines, namely on the a) INSTA data, b) NBS data and c) PointAvatar data. The close-ups on the right of each example highlight our method’s ability to capture details like teeth, wrinkles and reflections.
@@ 4.2 Ablations
In this section we ablate the HeadGaS components. First, we train a model that does not use learned features for blending, but rather a base of colors and opacities, and uses the expression weights to obtain the final color and opacity as a weighted average. We refer to this model as Ours w/o MLP. In addition, since an intuitive alternative for dynamic Gaussians would be to deform the points (rather than adapting color and opacity) we introduce Ours w/ $\Delta(\mu,R)$, which uses the learned feature basis (and a similar MLP) to rather shift the positions $\mu$ and rotations $R$. Additionally, we train a model that predicts color, opacity, as well as a $\mu$ and $R$ shift (Ours change all). Further, we run a model without the perceptual loss, i.e. Ours w/o $\mathcal{L}_{p}$, to ablate its effectiveness. Additionally, to validate the contribution of using the expression parameters as a weight for blending Gaussian features, rather than a simple condition to the MLP, we ablate a variant named Ours w/o blending. For fairness, we increase the capacity of this MLP until it plateaus. More details can be found in the supplement.
Table 2 reports the quantitative evaluation of our model components. We observe that using the expression parameters as a simple condition (Ours w/o blending) leads to noticeably inferior performance. In contrast to our model - which learns per-Gaussian dynamics via a feature basis - the per-subject MLP has to learn the face dynamics for all Gaussians at once, leading to poorer expression generalization, as can be also seen on Figure 5. Also, applying a transformation to the positions and rotations leads to worse results (Ours w/ $\Delta(\mu,R)$). We hypothesize this is because, in the context of 3DGS - which is relying on several heuristics - adding another dimension (in the form of spatial 3D motion) further complicates the already difficult optimization, resulting in geometrically inconsistent transformation (e.g. failure to preserve relative distances of points in the skin), especially for large motion. Figure 5 confirms these results and reveals floater artifacts and less accurate expressions. We also observe that allowing all parameters to change (Ours change all) increases the solution space and makes the heuristic-based 3DGS optimization more challenging, leading to blurrier results. Further, blending the explicit parameters directly (Ours w/o MLP) leads to a worse performance than our neural variant. Despite a tighter PSNR gap, we notice a drastic visual effect on the highly dynamic areas, relevant to the blending, as illustrated in Figure 5. Finally, we see that adding a perceptual loss term $\mathcal{L}_{p}$ leads to an improvement in most metrics. We refer the reader to the supplement for video comparisons, as well as ablation on additional aspects such as number of Gaussians and speed.
Table 2: Ablation of HeadGaS components on the INSTA dataset.
@@
Figure 5: Qualitative ablation on the INSTA dataset
@@ 4.3 Novel View Synthesis
In Figure 7 we render the avatars from multiple views, including the original test set camera (left) and two additional viewpoints (right). Thereby we render the same facial expression. We observe that our model can deliver expressions that are consistent across different views. Videos can be found in the supplement.
Figure 6: Rendering from various camera views for the same expressions
Figure 7: Cross-subject driving: rendering target subject with expression from the source subject
@@ 4.4 Cross-Subject Expression Driving
Figure 7 reports our cross-subject driving results, i.e. use the facial expression from another (ground truth) source subject to drive a target subject. Here we retain the original head pose of the target subject. As can be seen in the figure, our model is capable of transferring various expressions, e.g. talking, wink, surprise, across different subjects at a reasonable quality.
Figure 8: Limitations. Left: subjects only observed in neutral expression in non-frontal view. Right: Camera view that is far from the training observations.
@@ 5 Limitations and Ethical Consideration
HeadGaS is affected by failures of the head tracker, which can take the form of inaccurate facial expressions, or blurriness in case of pose inaccuracy in the training data (Figure 8). Further, as a data driven method, HeadGaS requires a reasonable coverage of expressions across different views. For instance, if a head changes pose only in neutral expression, and diverse expressions are observed only in frontal view, it would be difficult to capture a non-neutral expression from a side view (Figure 8, left). Finally, a downside of HeadGaS is the memory consumption originating from the feature bases ($B\times f_{dim}$ floats per Gaussian).
The use of personal data should be handled carefully, and follow local regulations. We note that face reanimation can potentially be used to generate fake content, and convey misinformation. We do not condone such practises, and believe that the community should work accordingly towards mitigating the risks.